Calculate the internal energy and enthalpy of 1kg of air occupying 0.05m
1
Expert's answer
2021-07-19T03:14:53-0400

We will assume that gas is ideal. So, the internal energy of air



"U=\frac{m}{M}C_VT"



For air "M=28.96 \cdot 10^{-3}kg/mol" and "C_V=\frac{i}{2}R=\frac{5}{2}\cdot 8.31\approx20.8\frac {J}{mol\cdot K}"



"U=\frac{m}{M}C_VT=\frac{1}{28.96 \cdot 10^{-3}}\cdot 20.8\cdot 293=210442J"



The enthalpy of air


"H=U+pV=210442+2\cdot 10^6\cdot 0.05=310442J"